# Safe Pareto Improvements for Expected Utility Maximizers in Program Games

### v4

@@ 4.2 Components of CSR
How can a player guarantee a payoff better than some miscoordination outcome, without losing the opportunity to bargain for their most-preferred outcome? Suppose player $i$ considers using a set-valued renegotiation function that, against some counterpart profile, achieves an outcome $\boldsymbol{a}$ that is worse for at least one player than their least-preferred efficient outcome (i.e., their PMM payoff). As we will argue in Theorem 3, under mild assumptions, player $i$ is no worse off also including their PMP of $\boldsymbol{a}$ in their renegotiation set. On the other hand, if $i$ thinks the selection function might not choose the best outcome for themselves in the agreement set, $i$ will not prefer to include outcomes strictly better for some player $j$ than those in their PMP of $\boldsymbol{a}$. For example, in Fig. 2, player 1’s set safely guarantees the PMM against players who also use sets of this form. But if player 1’s set in the right plot were any larger — and hence included outcomes worse for player 1 than player 1’s best possible outcome — the selection function might choose an outcome that is worse for player 1 than otherwise.
Notice that players’ ability to guarantee the PMM depends on conditional renegotiation sets, for the reasons discussed in Example 4.2: If player 1 added the outcome marked with the star in Fig. 3 to an unconditional renegotiation set, in an attempt to Pareto-improve to their PMM payoff, this could make them worse off against a player 2 with the renegotiation set shown in the right plot.
@@
Finally, here is the formal definition of CSR programs. For a set-valued renegotiation function $\mathbf{RN}^{i}$, we define the space of CSR programs $P_{i}^{\textnormal{RN}}(\mathbf{RN}^{i})$ as the space of programs with the structure of Algorithm 2, for some default $p_{i}^{\textnormal{def}}$. (Let $\mathcal{R}^{\textnormal{RN}}$ be the set of all set-valued renegotiation functions, and for each $i$ let $P_{i}^{\textnormal{RN}}=\bigcup_{\mathbf{RN}^{i}\in\mathcal{R}^{\textnormal{RN}}}P_{i}^{\textnormal{RN}}(\mathbf{RN}^{i})$, i.e., the space of all CSR programs.) The selection function $\mathbf{sel}$ is given to the players (and we suppress dependence of $P_{i}^{\textnormal{RN}}$ on $\mathbf{sel}$ for simplicity).
@@
1:Counterpart program profile $\textbf{p}_{-i}$
2:if $\textbf{p}_{-i}\in\bigtimes_{j\neq i}P_{j}^{\textnormal{RN}}(\mathbf{RN}^{j})$ for some $(\mathbf{RN}^{j})_{j\neq i}\in\bigtimes_{j\neq i}\mathcal{R}^{\textnormal{RN}}$ then
@@
4:     ${I}\leftarrow\bigcap_{j=1}^{{n}}\mathbf{RN}^{j}(\mathbf{RN}^{-j},\widehat{\boldsymbol{a}})$triangleright Agreement set
@@
9:     return $p_{i}^{\textnormal{def}}(\textbf{p}_{-i})$
4.3 Guaranteeing PMM payoffs using CSR
Similar to the assumption in Section 3.4, suppose players always include more outcomes in their renegotiation sets if their expected utility is unchanged. Then, we will show in Theorem 3 that under mild assumptions on players’ beliefs, for any program that does not guarantee a player at least their PMM payoff, there is a corresponding CSR program the player prefers that does guarantee the PMM payoff. We prove this result by constructing programs identical to the programs players would otherwise use, except that these new programs’ renegotiation sets for each outcome include their PMP of the outcome they would have otherwise achieved. For a program $p_{i}$, we call this modified program the PMP-extension of $p_{i}$ (Definition 8).
This result requires two assumptions on players’ beliefs and the structure of programs used in subjective equilibrium (Assumptions 9i and 9ii), which are analogous to Assumption 5 of Proposition 2:
  1. 1.
@@ 4.3 Guaranteeing PMM payoffs using CSR
  2. 2.
    Informally, Assumption 9ii says that players believe that, with probability 1: If a CSR program is modified only by adding PMP points to its renegotiation set, the only changes the counterparts would prefer to make are those that also add PMP points. The intuition for this assumption is that the PMP-extension, by definition, doesn’t add any points that make counterparts strictly better off than some possible renegotiation outcome while making the focal player worse off. So, the counterparts don’t have incentives to make changes to their renegotiation sets that would make the focal player worse off. (This argument wouldn’t work if player $i$ also added outcomes that are better for some counterpart $j$ than their PMP-extension. This is because, as noted in the previous section, $j$ would then have an incentive to exclude  $i$’s most-preferred outcome from $j$’s renegotiation set.)
@@
Setup for Theorem 3.
Consider any set-valued renegotiation functions $\mathbf{RN}^{i},\mathbf{RN}^{-i}$, CSR programs $\textbf{p}\in\bigtimes_{j=1}^{{n}}P_{j}^{\textnormal{RN}}(\mathbf{RN}^{j})$ with selection function $\mathbf{sel}$, and $\boldsymbol{a}\in\mathcal{A}$. Denote the agreement set as $\textbf{AS}(\mathbf{RN}^{i},\mathbf{RN}^{-i},\boldsymbol{a})=\bigcap_{j=1}^{{n}}\mathbf{RN}^{j}(\mathbf{RN}^{-j},\boldsymbol{a})$. Let the outcome of renegotiation from point $\boldsymbol{a}$ using CSR programs p, if any, be
$\mathbf{ro}(\textbf{p},\boldsymbol{a})$ =$=\begin{cases}\mathbf{sel}(\textbf{AS}(\mathbf{RN}^{i},\mathbf{RN}^{-i},\boldsymbol{a})),&\text{if }\textbf{AS}(\mathbf{RN}^{i},\mathbf{RN}^{-i},\boldsymbol{a})\neq\emptyset,\\ \boldsymbol{a},&\text{else.}\end{cases}$
Now define the following:

### v5

@@ 4.2.3. Renegotiation sets that guarantee the PMM
    On the other hand, if $i$ thinks the selection function might not choose their optimal outcome in the agreement set, $i$ will not prefer to include outcomes strictly better for player $j$ than those in $\mathbf{PMP}_{i}(\boldsymbol{a})$. For example, in Fig. 2, $\mathbf{RN}^{1}(\mathbf{RN}^{2},(0,0))$ includes all outcomes Pareto-worse than $\mathbf{PMP}_{1}((0,0))$. This set safely guarantees the PMM against a player who also uses a set of this form, and gives player 1 their best possible outcome against the $\mathbf{RN}^{2}$ in the right plot. But if $\mathbf{RN}^{1}(\mathbf{RN}^{2},(0,0))$ in the right plot included additional outcomes, which would be worse for player 1 than player 1’s best possible outcome, the selection function might choose an outcome that is worse for player 1 than otherwise. (This is why, when we construct strategies for the proof of Theorem 9, we add the entire PMP even though it is sufficient to only add the point in the PMP that minimizes the player’s payoff.)
@@
Finally, here is the formal definition of CSR programs. For a set-valued renegotiation function $\mathbf{RN}^{i}$, we define the space of CSR programs $P_{i}^{\textnormal{RN}}(\mathbf{RN}^{i})$ as the space of programs with the structure of Algorithm 2, for some default $p_{i}^{\textnormal{def}}$. (Let $\mathcal{R}^{\textnormal{RN}}$ be the set of all set-valued renegotiation functions, and for each $i$ let the space of all CSR programs be $P_{i}^{\textnormal{RN}}=\bigcup_{\mathbf{RN}^{i}\in\mathcal{R}^{\textnormal{RN}}}P_{i}^{\textnormal{RN}}(\mathbf{RN}^{i})$. Let $\mathcal{P}^{\text{RN}}=P_{1}^{\textnormal{RN}}\times P_{2}^{\textnormal{RN}}$.) The selection function $\mathbf{sel}$ is given to the players (and we suppress dependence of $P_{i}^{\textnormal{RN}}$ on $\mathbf{sel}$ for simplicity).
@@
1:Counterpart program $p_{j}$
2:if $p_{j}\in P_{j}^{\textnormal{RN}}(\mathbf{RN}^{j})$ for some $\mathbf{RN}^{j}\in\mathcal{R}^{\textnormal{RN}}$ then
@@
4:     ${I}\leftarrow\mathbf{RN}^{1}(\mathbf{RN}^{2},\widehat{\boldsymbol{a}})\cap\mathbf{RN}^{2}(\mathbf{RN}^{1},\widehat{\boldsymbol{a}})$triangleright Agreement set
@@
9:     return $p_{i}^{\textnormal{def}}(p_{j})$
4.3. Guaranteeing PMM Payoffs Using CSR
Similar to the assumption in Section 3.3, suppose players always include more outcomes in their renegotiation sets if their expected utility is unchanged. So in particular, to show that in subjective equilibrium players use programs that guarantee at least the PMM, it will suffice to show that they weakly prefer these programs.
Then, we will show in Theorem 9 that under mild assumptions on players’ beliefs, for any program that does not guarantee a player at least their PMM payoff, there is a corresponding CSR program the player prefers that does guarantee their PMM payoff. We prove this result by constructing programs identical to the programs players would otherwise use, except that these new programs’ renegotiation sets for each outcome include their PMP of the outcome they would have otherwise achieved. For a program $p_{i}$, we call this modified program the PMP-extension of $p_{i}$ (Definition 8).
This result requires two assumptions on players’ beliefs and the structure of programs used in subjective equilibrium (Assumptions 7i and 7ii), analogous to Assumption 1 of Proposition 5:
  1. (1)
@@ 4.3. Guaranteeing PMM Payoffs Using CSR
  2. (2)
    Informally, Assumption 7ii says that players believe that, with probability 1: If a CSR program is modified only by adding PMP points to its renegotiation set, the only changes the counterparts would prefer to make are those that also add PMP points. The intuition for this assumption is: For any possible default renegotiation outcome, the PMP-extension, by definition, doesn’t add any points that make the counterpart strictly better off than that outcome while making the focal player worse off (see Fig. 4). So, similar to Assumption 7i, the counterpart doesn’t have an incentive to make changes to their renegotiation set that would make the focal player worse off. (This argument wouldn’t work if player $i$ also added outcomes that are better for $j$ than their PMP-extension. This is because, as noted in the previous section, $j$ would then have an incentive to exclude  $i$’s most-preferred outcome from $j$’s renegotiation set.)
@@
Definition 1.
For any $p_{i}\in P_{i}^{\textnormal{RN}}(\mathbf{RN}^{i})$ for some $\mathbf{RN}^{i}$, the PMP-extension $\widetilde{p_{i}}\in P_{i}^{\textnormal{RN}}({\widetilde{\mathbf{RN}}}^{i})$ is the program identical to $p_{i}$ except: for all $p_{j}\in P_{j}^{\textnormal{RN}}(\mathbf{RN}^{j})$ for some $\mathbf{RN}^{j}$, writing ${\widetilde{\textbf{p}}^{i}}=(\widetilde{p_{i}},p_{j})$, we have
${\widetilde{\mathbf{RN}}}^{i}(\mathbf{RN}^{j},\boldsymbol{a}({\widetilde{\textbf{p}}^{i^{\textnormal{def}}}}))$ $=\mathbf{RN}^{i}(\mathbf{RN}^{j},\boldsymbol{a}({\widetilde{\textbf{p}}^{i^{\textnormal{def}}}}))\cup\mathbf{PMP}_{i}(\boldsymbol{a}(\textbf{p})).$
Assumption 2.
  1. (i)
    Suppose either $p_{i}$ is in a subjective equilibrium of $G(\mathcal{P}\cup\mathcal{P}^{\text{RN}})$, or $p_{i}$ is in the support of $\beta_{j}$. Suppose $p_{i}\notin P_{i}^{\textnormal{RN}}$. Then for any $p_{j}\in P_{j}^{\textnormal{RN}}$, we have $p_{i}(p_{j})=p_{i}(p_{j}^{\textnormal{def}})$.
  2. (ii)
    Let $p_{j}\in P_{j}^{\textnormal{RN}}(\mathbf{RN}^{j})$ be in the support of $\beta_{i}$, and take any $p_{i}\in P_{i}^{\textnormal{RN}}(\mathbf{RN}^{i})$ with PMP-extension $\widetilde{p_{i}}$. For all $\boldsymbol{a}$, we have that $\mathbf{RN}^{j}({\widetilde{\mathbf{RN}}}^{i},\boldsymbol{a})=\mathbf{RN}^{j}(\mathbf{RN}^{i},\boldsymbol{a})\cup{V}$ for some ${V}\subseteq\mathbf{PMP}_{i}(\boldsymbol{a}(\textbf{p}))$.
$x$$y$
$\boldsymbol{a}({\textbf{p}}^{\textnormal{def}})$
$\boldsymbol{a}(\textbf{p})$
$\mathbf{PMP}_{1}(\boldsymbol{a}(\textbf{p}))$
Figure 4. Illustration of the argument for Theorem 9. By default, the renegotiation outcome is the black circle, $\boldsymbol{a}(\textbf{p})$. Player 1 considers whether to add to their renegotiation set $\mathbf{RN}^{1}(\mathbf{RN}^{2},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))$ the black striped segment $\mathbf{PMP}_{1}(\boldsymbol{a}(\textbf{p}))$. Player 1 is certain that player 2 would not change their set $\mathbf{RN}^{2}(\mathbf{RN}^{1},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))$ in response to this addition in a way that would make player 1 worse off (Assumption 7). This is because the only change player 1 has made is to add outcomes that make both players weakly better off than $\boldsymbol{a}(\textbf{p})$ and do not make player 2 strictly better off.
[…]
The default renegotiation outcome a(p) is marked with a black circle. A black striped segment shows Player 1’s PMP of this outcome, PMP₁(a(p)), which includes points that make both players weakly better off than a(p) while not making Player 2 strictly better off. The shaded region represents Player 1’s renegotiation set RN¹(RN², a(p^def)).
Theorem 9.
Let $G(\mathcal{P})$ be a program game, and $\mathbf{sel}$ be any transitive selection function. Suppose the action sets of $G$ are continuous, so that for any $\boldsymbol{a}\in\mathcal{A}$, player $i$’s PMP of that action profile $\mathbf{PMP}_{i}(\boldsymbol{a})$ is nonempty. Let $\boldsymbol{\beta}$ be any belief profile satisfying the assumption that players are (i) certain that CSR won’t be punished and (ii) certain that PMP-extension won’t be punished (Assumption 7).
Then, for any subjective equilibrium $(\textbf{p},\boldsymbol{\beta})$ of $G(\mathcal{P}\cup\mathcal{P}^{\text{RN}})$ where $U_{i}(\textbf{p})<u_{i}^{\textnormal{PMM}}$ for some $i$, there exists $\textbf{p}^{\prime}$ such that:
  1. (1)
    For all $i$, $p_{i}^{\prime}$ is the PMP-extension of $p_{i}$.
  2. (2)
    $\textbf{U}(\textbf{p}^{\prime})\succeq\textbf{u}^{\textnormal{PMM}}$.
  3. (3)
    $(\textbf{p}^{\prime},\boldsymbol{\beta})$ is a subjective equilibrium of $G(\mathcal{P}\cup\mathcal{P}^{\text{RN}})$.
Proof Sketch.
Assumption 7i implies that players always use CSR programs. Consider any renegotiation outcome $\boldsymbol{a}$ worse for some player than the PMM, which is achieved by player $i$’s “old” program against some counterpart. By Assumption 7ii, player $j$ doesn’t punish $i$ for adding their PMP of that outcome, $\mathbf{PMP}_{i}(\boldsymbol{a})$, to their renegotiation set (in their “new” program). So the renegotiation outcome of the new program against $j$ is only different from that of the old program if $j$ is also willing to renegotiate to some outcome in $\mathbf{PMP}_{i}(\boldsymbol{a})$. But in that case, because the selection function is transitive, the new renegotiation outcome is no worse for $i$ than under the old program. Therefore, each player always prefers to replace a given program with its PMP-extension, and when all players use PMP-extended programs, the Pareto frontier of their agreement set only includes outcomes guaranteeing each player their PMM payoff. $\square$
Remark: Notice that the argument above does not require that players refrain from using programs that implement other kinds of SPIs, besides PMP-extensions. First, the PMP-extension can be constructed from any default program, including, e.g., a CSR program whose renegotiation set is only extended to include the player’s best outcome, not their PMP (call this a “self-favoring extension”). And if a player’s final choice of program is their self-favoring extension, they are still incentivized to use the PMP-extension within their default program.
Second, while it is true that an analogous argument to the proof of Theorem 9 could show that a player is weakly better off ex ante using a self-favoring extension than not extending their renegotiation set at all, this does not undermine our argument. This is because, as we claimed at the start of this section, it is reasonable to assume that among programs with equal expected utility, each player prefers to also include their PMP. But wouldn’t the player also prefer an even larger renegotiation set that includes outcomes that Pareto-dominate the PMM as well? No, because those outcomes will be worse for that player and better for their counterpart than the player’s most-preferred outcome, such that the counterpart would have an incentive to make the player worse off (i.e., it’s plausible that Assumption 7ii would be violated).
We can now formalize the claim that CSR is an SPI that partially solves SPI selection: The mapping from programs p to instances of Algorithm 2 withp as defaults, for any profile $(\mathbf{RN}^{1},\mathbf{RN}^{2})$ used in subjective equilibrium under the assumptions of Theorem 9, is an SPI that guarantees players their PMM payoffs.
Proposition 10.
For $i=1,2$, for some selection function $\mathbf{sel}$, define $f_{i}^{\mathbf{RN}^{i}}:P_{i}\rightarrow P_{i}^{\textnormal{RN}}(\mathbf{RN}^{i})$ such that, for each $p_{i}\in P_{i}$, $f_{i}^{\mathbf{RN}^{i}}(p_{i})$ is of the form given in Algorithm 2 with ${f_{i}^{\mathbf{RN}^{i}}(p_{i})}^{\textnormal{def}}=p_{i}$. Then, under the assumptions of Theorem 9, for any $(\mathbf{RN}^{1},\mathbf{RN}^{2}),{\textbf{p}}^{\textnormal{def}}$ such that for all $\mathbf{RN}^{j}$, $\mathbf{PMP}_{i}(\boldsymbol{a}(\textbf{p}))\subseteq\mathbf{RN}^{i}(\mathbf{RN}^{j},\boldsymbol{a}(\textbf{p}))$:
  1. (1)
    The function $\textbf{f}^{\mathbf{RN}}:\textbf{p}\mapsto(f_{1}^{\mathbf{RN}^{1}}(p_{1}),f_{2}^{\mathbf{RN}^{2}}(p_{2}))$ is an SPI.
  2. (2)
Proof.
This follows immediately from the argument used to prove Theorem 9. ∎
In Appendix D, we show that players are not always incentivized to use SPIs that strictly improve on the PMM.
Table 2. Key notation
Symbol Description (page introduced)
$\boldsymbol{a}(\textbf{p})$ action profile in the base game played by players
with the given program profile (2)
$\mathbf{rn}^{i}$ renegotiation function for player $i$ (maps an action
profile to a Pareto-improved action profile) (3)
$\mathbf{RN}^{i}$ set-valued renegotiation function for player $i$
(maps $j$’s set-valued renegotiation function and
an action profile to a set of Pareto-improved
action profiles) (5)
$\mathcal{R},\mathcal{R}^{\textnormal{RN}}$ sets of all renegotiation functions and set-valued
renegotiation functions, respectively (3, 6)
$P_{i}^{\textnormal{rn}}(\mathbf{rn}^{i})$ set of renegotiation programs (Algorithm 1) that
use the renegotiation function $\mathbf{rn}^{i}$ (3)
$P_{i}^{\textnormal{RN}}(\mathbf{RN}^{i})$ set of conditional set-valued renegotiation
programs (Algorithm 2) that use the set-valued
renegotiation function $\mathbf{RN}^{i}$ (6)
$p_{i}^{\textnormal{def}}$ default program for a program $p_{i}$ in $P_{i}^{\textnormal{rn}}$ or $P_{i}^{\textnormal{RN}}$ (3)
$\mathbf{sel}$ selection function (maps a set of action profiles to
an action profile that is efficient within that set) (5)
$\textbf{u}^{\textnormal{PMM}}$ Pareto meet minimum (5)
$\mathbf{PMP}_{i}$ Pareto meet projection for player $i$ (maps an
action profile to a particular set of Pareto-improved
action profiles) (5)
5. Discussion
Using renegotiation to construct SPIs in program games is a rich and novel area, with many directions to explore. To name a few:
  • Which plausible conditions would violate our assumptions about players’ beliefs used for the PMM guarantee?
  • What do unilateral SPIs (Oesterheld and Conitzer, 2022) look like in this setting?
  • When are SPIs used in sequential, rather than simultaneous-move, settings? In particular, in sequential settings, the first-moving player’s decision whether to use a renegotiation program could signal private information to the second-moving player.
  • We have assumed complete information about payoffs; using ideas from DiGiovanni and Clifton (2023)’s framework for program games in the presence of private information, it should also be possible to construct SPIs in incomplete information settings.
  • How can this theory inform real-world AI system design?
{acks}
For valuable comments and discussions, we thank our anonymous referees, Caspar Oesterheld, Lukas Finnveden, Vojtech Kovarik, and Alex Kastner.
Appendix A ${n}$-Player Notation and Proof of Proposition 5
We extend the 2-player formalism in the main text to the ${n}$-player case as follows (all other extensions from the 2-player case to the ${n}$-player case are straightforward):
  • In a program game with program space $\mathcal{P}$, each player $i$’s beliefs $\beta_{i}$ are a distribution supported on ×$\bigtimes_{j\neq i}P_{j}$.
  • A profile of programs, renegotiation functions, etc. with the subscript $-i$ denotes the profile with the $i$th entry excluded. E.g., $\textbf{p}_{-i}=(p_{j})_{j\neq i}$.
  • $\mathbf{PMP}_{i}(\boldsymbol{a})$ is the set of action profiles $\tilde{\boldsymbol{a}}$ such that $u_{i}(\tilde{\boldsymbol{a}})\geq\max\{u_{i}^{\textnormal{PMM}},u_{i}(\boldsymbol{a})\}$ and $u_{j}(\tilde{\boldsymbol{a}})=\max\{u_{j}^{\textnormal{PMM}},u_{j}(\boldsymbol{a})\}$ for all $j\neq i$.
  • For a profile $\textbf{p}\in\bigtimes_{j=1}^{n}P_{j}^{\textnormal{RN}}(\mathbf{RN}^{j})$, let $\mathbf{RN}^{-i}=(\mathbf{RN}^{j})_{j\neq i}$.
The statement of Assumption 1 in the general ${n}$-player case is:
Assumption 4.
  1. (1)
  2. (2)
In this context, let $\mathcal{P}^{\text{rn}}=\bigtimes_{i=1}^{{n}}P_{i}^{\textnormal{rn}}$. Then:
See 5
Proof.
Let $(\textbf{p}^{*},\boldsymbol{\beta})$ be a subjective equilibrium of $G(\mathcal{P}\cup\mathcal{P}^{\text{rn}})$, where for some $i$, $p_{i}^{*}\notin P_{i}^{\textnormal{rn}}$; and let $\mathbf{rn}$ be an arbitrary renegotiation function. Let $\textbf{p}_{-i}$ be in the support of $\beta_{i}$. We will show that against any such $\textbf{p}_{-i}$, the new program $f_{i}^{*}(p_{i}^{*})\in P_{i}^{\textnormal{rn}}(\mathbf{rn})$ is weakly better for player $i$ than $p_{i}^{*}$ (hence $f_{i}^{*}(p_{i}^{*})$ is better in expectation).
Now, let $\textbf{p}^{\prime}$ be the program profile such that:
  • $p_{j}^{\prime}=p_{j}^{*}$ if $p_{j}^{*}\in P_{j}^{\textnormal{rn}}(\mathbf{rn}^{j})$ for some $\mathbf{rn}^{j}$
  • $p_{j}^{\prime}=f_{j}^{*}(p_{j}^{*})$ otherwise.
Clearly, then, (1) holds by construction, and each $p_{j}^{\prime}\in P_{j}^{\textnormal{rn}}$. And since the above argument holds for any $j$, we have
$f_{j}^{*}(p_{j}^{*})\in\operatorname*{arg\,max}_{p_{j}\in P_{j}\cup P_{j}^{\textnormal{rn}}}\mathbb{E}_{\textbf{p}_{-j}\sim\beta_{j}}U_{j}(\textbf{p}),$
i.e., (2) $(\textbf{p}^{\prime},\boldsymbol{\beta})$ is a subjective equilibrium. ∎
Appendix B Proof of Claim that Coordination on the Selection Function Does Not Change Players’ Payoffs
Appendix C Proof of Theorem 9
Denote the agreement set as $\textbf{AS}(\mathbf{RN}^{i},\mathbf{RN}^{-i},\boldsymbol{a})=\bigcap_{j=1}^{n}\mathbf{RN}^{j}(\mathbf{RN}^{-j},\boldsymbol{a})$.
In the ${n}$-player setting, define: